# {strip}[Uncaptioned image]

### v1

@@ Abstract.
Research on diffusion model-based video generation has advanced rapidly. However, limitations in object fidelity and generation length hinder its practical applications. Additionally, specific domains like dynamic wallpapers require seamless looping, where the first and last frames of the video match seamlessly. To address these challenges, this paper proposes LoopAnimate, a novel method for generating videos with consistent start and end frames. To enhance object fidelity, we introduce a framework that decouples multi-level image appearance and textual semantic information. Building upon an image-to-image diffusion model, our approach incorporates both pixel-level and feature-level information from the input image, injecting image appearance and textual semantic embeddings at different layers of the diffusion model. Existing UNet-based video generation methods require to input the entire videos during training to encode temporal and positional information at once. However, due to limitations in GPU memory, the number of frames is typically restricted to 16. To address this, this paper proposes a three-stage training strategy with progressively increasing frame numbers and reducing fine-tuning modules. Additionally, we introduce the Temporal Enhanced Motion Module (TEMM) to extend the capacity for encoding temporal and positional information up to 36 frames. The proposed LoopAnimate, which for the first time extends the single-pass generation length of UNet-based video generation models to 35 frames while maintaining high-quality video generation. Experiments demonstrate that LoopAnimate achieves state-of-the-art performance in both objective metrics, such as fidelity and temporal consistency, and subjective evaluation results.
Diffusion, Image to video generation, Loopable, Long video
@@ 1. Introduction
Diffusion model based video generation  (liu2024sora, ; guo2023animatediff, ; chen2023livephoto, ; blattmann2023stable, ; wang2023videocomposer, ; tuneAvideo, ; magicavatar, ) is a recent research hotspot filled with challenges. Existing approaches can be categorized into two main architectures: DiT-based  (peebles2023scalable, ) and UNet-based  (ronneberger2015unet, ). DiT-based methods, such as SORA (liu2024sora, ) and Latte (ma2024latte, ), utilize Diffusion Transformers (DiT) (peebles2023scalable, ) as their backbone. While these methods demonstrate potential for generating longer and higher-quality videos, they also demand significant computational resources and longer inference times. UNet-based methods, including Animatediff (guo2023animatediff, ), Make Pixel Dance (zeng2023make, ), PIA(zhang2023pia, ), and LivePhoto (chen2023livephoto, ), represent the other prevalent approaches. These models typically require inputting the entire video sequence during training and employ motion modules (guo2023animatediff, ) for temporal and positional encoding. However, limited by GPU memory constraints, most current works are restricted to generating 16 frames. Focusing on practical applications, this paper proposes LoopAnimate, a novel image-to-video generation method specifically designed for dynamic wallpaper. LoopAnimate addresses the challenge of generating longer video sequences by enabling seamless transitions between the first and last frames, achieving a total output of 35 frames at once.
To address the challenge of object fidelity in image-to-video generation tasks, we propose the Multilevel Image representation and Textual semantics Decoupling Framework (MITDF). For long video generation, we introduce a Three-Stage Training Strategy to progressively increase frames while reducing fine-tuning modules, along with a TEMM that extends temporal positional encoding capabilities to 36 frames. Within MITDF, we utilize a pretrained image-to-image diffusion model (sd-image-variations-diffusers, ) as the foundation. Input image information is incorporated into the generation process at both pixel-level and feature-level through conditioning and image embedding, respectively. Image representation and textual semantic embedding are injected into different stages of the diffusion model via cross-attention. Experiments demonstrate that excessive image embedding information will limit the motion dynamics of the generated videos, while the middle block contributes most significantly. Consequently, image representation embedding is injected only during the down-sampling process of the diffusion model, and textual semantic embedding is injected solely into the middle and up-sampling blocks. Additionally, we observe that this architecture effectively mitigates the influence of WebVid10M training data  (Bain21, ) watermarks during inference.
@@
  • Long Video Generation. We present a Three-Stage Training Strategy that progressively increases the number of generated frames in each stage while reducing the fine-tuning modules. Combined with the TEMM, which extends the temporal encoding length to 36 frames, our approach enables the generation of 35-frame videos in a single pass.
@@ 2.2. Long Video Generation
  • Stable Video Diffusion  (blattmann2023stable, ) presents a latent video diffusion model for high-resolution, state-of-the-art image-to-video generation. And SVD-XT can generate good quality image-conditioned videos with 25 frames in 2 to 8 steps with 576x1024 resolutions.
@@ 3. Method
Figure 2. Illustration of LoopAnimate.
@@
Multi-level image representation leverages information from both the pixel-level and feature-level of the input image. As illustrated in Figure 2, during training, the first and last frames (marked as $F(0)$ and $F(-1)$) of the video are used as one of the conditions to the UNet in the form of Static and Dynamic Stacking Latent Code (SDSLC), enabling pixel-level information injection. Additionally, image embedding is extracted from the input image and cross-attention is performed during the down-sampling process, encoding the image representation information into the latent space and achieving feature-level information injection.
Appearance-semantic decoupling involves injecting the image representation embedding and text semantic embedding at different layers of the denoising network. Inspired by  (basu2023localizing, ), which suggests that different layers of the denoising network in text-to-image tasks have varying influences on style, objects, color, and motion, we conducted ablation experiments in Sec. 4.2.1. Similar to text-to-image tasks, we found that the middle block has the most significant impact on motion. We further discovered that excessive insertion of image embedding will restrict the motion dynamics of the generated video. Based on these findings, we finalized the network architecture as shown in Figure 2. Additionally, we increased the temporal position encoding length in the temporal attention module within the motion module, naming it the Temporal Enhanced Motion Module. We were pleasantly surprised to find that this structure effectively avoids the influence of watermarks present in the WebVid10M training data during inference.
In this section, we introduce the (1) fundamentals of diffusion models for image and video generation as a foundation. Subsequently, the proposed method is presented in detail, following the order of the (2) Asymmetric Loop Sampling Strategy for training data, the (3) specially designed condition initialization method and the (4) MITDF, and the (5) three-stage training strategy.
@@ 3.3.3. Fine-Grained Semantic Mask
During the inference stage, the object’s mask $M_{0}$ is obtained from the input image. We then create a tensor $Z_{M}$ by replicating one padding to all frames except for the first and last frames ${Z}_{M}={[{M}_{0},OnePAD,{M}_{0}]}$. Finally, $Z_{t}$, $Z_{SD}$, and $Z_{M}$ are concatenated along the channel dimension to form the final input condition ${Z}_{C}$ for the UNet, denoted as Equation (11), where ${{Z}_{C}}\in\mathcal{R}^{n*9*(2f-1))*h*w}$,
@@
3.4. Image representation and Textual semantics Decoupling Framework
Object fidelity is a crucial aspect of image-to-video generation. To address this challenge, we propose a framework built upon a pretrained image-to-image diffusion model which initially utilizes image embedding as the guidance, and we enable the capability of decoupling multi-level image appearances and textual semantics. This decoupling involves injecting image appearance embedding and text semantic embedding at different layers of the denoising network. Multi-level image representation refers to incorporating pixel-level representation information of the input image during condition initialization, see Equation (8). And extracte image embedding from the input image to perform cross-attention to inject the feature-level information into the latent space. Our ablation experiments reveal that, similar to text-to-image tasks (rombach2022highresolution, ; pixart-alpha, ; Imagen, ; rapheal, ; qu2023layoutllm, ; sun2023sgdiff, ), text semantics in image-to-video tasks exert the most significant influence in the middle block. Through limited ablation experiments in Sec. 4.2.1, we ultimately determined the optimal structure representation to be Equation (12) and Equation (13).
@@ 3.4. Image representation and Textual semantics Decoupling Framework
Similar to Equation (14), as shown in Equation (15), ${Q}_{Z}^{M,U}$ represents the query extracted from the latent code ${Z}$ of the middle and up-sample layers, while ${K}_{I}$ and $V_{I}$ denote the key and value obtained from the image embedding, respectively.
@@
${{W_{q}},{W_{k}},{W_{v}}}$ are the weight matrices of the trainable linear projection layers in down-sample blocks, and ${{W_{q}^{{}^{\prime}}},{W_{k}^{{}^{\prime}}},{W_{v}^{{}^{\prime}}}}$ are in middle and up-sample blocks. ${{c_{t}},{c_{i}}}$ are text embedding and image embedding separately.
@@ 3.5. Three-Stage Training Strategy
Current UNet-based video generation models require the input of an entire video for temporal and positional information encoding. However, limited by hardware memory, most existing video generation works can only generate $16$ frames at a time. To address this issue, we propose a three-stage training strategy which progressively increases the number of generated frames while reducing the fine-tuning modules in each stage. Additionally, we introduce the TEMM, which extends the temporal and positional information encoding length to $36$ frames.
@@
Stage2. When the input image contains a salient object, maintaining object consistency and action continuity in the generated results becomes crucial, increasing the difficulty of this specific type of video generation task. To tackle this challenge, we leverage SOD  (liu2024lightweight, ) to determine whether the first frame of the video contains the same object as described in the caption. To identify whether the object exists within a video, we employed optical flow  (teed2020raft, ). This process allows us to curate the Sailent dataset, which contains 94,686 high-quality video segments. In the second stage, we train conv in and TEMM with $f=11$ ($21$ frames). Since the change of $f$ alters the encoding length of the positional embedding, TEMM needs to be fine-tuned. Additionally, fine-tuning conv in increases negligible computational cost but can accelerate convergence.
@@ 4. Experiments
Stage 2. The model is fine-tuned on a processed dataset of 94,686 high-quality videos containing salient objects (Sailent Dataset). The conv in layer and TEMM are fine-tuned with $f=11$, enabling the generation of 21 frame videos. The learning rate is fixed at $56.e-$ 5, and the fine-tuning process runs for 100k iterations.
Stage 3. The model is further fine-tuned on the Sailent Dataset, specifically focusing on conv in and the Q and V components of the temporal attention module within the TEMM. We increase $f$ to 18, enabling the generation of 35-frame videos. The learning rate is set to $52.e-$ 5, and the fine-tuning process runs for another 100k iterations. Table 1 shows the number of parameters fine-tuned in each stage. In Stage 3, only the conv in layer, TEMM.Q, and TEMM.V are fine-tuned, totaling $303.2M$ parameters. The number of frames selected for the three-stage training process is flexible. However, an excessively large number of frames in the first stage can prolong the training time. The temporal positional encoding length in TEMM can be extended, as long as it doesn’t exceed the GPU memory limit during the third stage fine-tuning. In our tests, using an A100 GPU with 80GB memory to run a single batch without any engineering optimization, the memory limit was reached when the third stage fine-tuning length was around 35 frames.
@@ 4.1. Comparison Results
We benchmarked our LoopAnimate algorithm against state-of-the-art image-to-video methods using objective metrics and subjective visual comparisons. We selected three open-source I2V methods, I2VGen-XL  (zhang2023i2vgenxl, ), PIA (zhang2023pia, ), and SVD-XT  (blattmann2023stable, ), as our baseline for comparison. We conducted experiments on a test set consisting of 99 images, and the results are presented in Table 2. In Stage 3, only the conv in, TEMM.Q, and TEMM.V modules were fine-tuned, achieved the best results across all metrics except Loop-C, while still being able to generate 35 video frames at once. The superior performance in CLIP-I and ${MSE}_{F_{0}}$ metrics indicates that our proposed method generates videos with high object fidelity, supporting the second point in our contributions regarding the effectiveness of the multi-level image appearance and textual semantic information decoupling framework. Furthermore, the three-stage fine-tuning strategy successfully extends the generation length to 35 frames while maintaining video quality, validating the third point in our contributions.
@@
I2VGen-XL (zhang2023i2vgenxl, ) 0.882 64.0 0.972 0.740 0.812 32
@@ 4.1.1. Visualization Results
Objective metrics fail to accurately assess the subjective visual quality of generated videos. Therefore, we opted for a subjective visual comparison experiment, employing several currently available methods: the open-source I2VGen-XL (zhang2023i2vgenxl, ), PIA (zhang2023pia, ), SVD-XT (blattmann2023stable, ) and Dynamicrafter (xing2023dynamicrafter, ), alongside the commercially available Runway Gen2  (gen2, ) and Pika1.0  (pikalab, ). The comparison results, as illustrated in Fig. 3, demonstrate that our proposed LoopAnimate method exhibits superior performance in both object fidelity and motion quality, particularly in the challenging task of generating real human portraits.
Furthermore, we conducted a user study involving 19 participants. Considering the extensive task volume, we selected 20 videos from a pool of 99 (as mentioned in Sec.  4.1) and evaluated the subjective quality based on two aspects: subject consistency ${Q}_{S}$ and motion quality ${Q}_{M}$. The scores ranged from 1 to 3, with 3 being the highest. The statistical results are presented in Table 3. Our LoopAnimate method achieved the highest scores for both ${Q}_{S}$ and ${Q}_{M}$. It should be noted that SVD-XT  (blattmann2023stable, ) lacks text input. Consequently, its ${Q}_{M}$ score is significantly lower due to the evaluation’s consideration of consistency with the provided textual prompts during the subjective assessment.

### v2

@@ Abstract.
Research on diffusion model-based video generation has advanced rapidly. However, limitations in object fidelity and generation length hinder its practical applications. Additionally, specific domains like dynamic wallpapers require seamless looping, where the first and last frames of the video match seamlessly. To address these challenges, this paper proposes LoopAnimate, a novel method for generating videos with consistent start and end frames. To enhance object fidelity, we introduce a framework that decouples multi-level image appearance and textual semantic information. Building upon an image-to-image diffusion model, our approach incorporates both pixel-level and feature-level information from the input image, through injecting image appearance and textual semantic embeddings at specific layers of the diffusion model. Existing UNet-based video generation methods use the entire videos as input during training to encode temporal and positional information at once. However, due to limitations in GPU memory, the number of frames is typically restricted to 16. To break this limit, this paper proposes a three-stage training strategy that progressively increases frame numbers and reduces fine-tuning modules. Additionally, we introduce the Temporal Enhanced Motion Module (TEMM) to extend the capacity for encoding temporal and positional information up to 36 frames. The proposed LoopAnimate extends generated video length of UNet-based video generation models to 35 frames while maintaining high-quality video generation. Experiments demonstrate that LoopAnimate achieves state-of-the-art performance in both objective metrics (e.g. fidelity and temporal consistency) and subjective evaluation results.
Diffusion, Image-to-video generation, Loopable video, Long video
@@ 1. Introduction
Diffusion model based video generation  (liu2024sora, ; guo2023animatediff, ; chen2023livephoto, ; blattmann2023stable, ; wang2023videocomposer, ; tuneAvideo, ; magicavatar, ) is a recent research hotspot filled with challenges. Existing approaches can be categorized into two main architectures: DiT-based  (peebles2023scalable, ) and UNet-based  (ronneberger2015unet, ). DiT-based methods, such as SORA (liu2024sora, ) and Latte (ma2024latte, ), utilize Diffusion Transformers (DiT) (peebles2023scalable, ) as their backbone. While these methods demonstrate potential for generating longer and higher-quality videos, they also demand significant computational resources and longer inference times. UNet-based methods, including Animatediff (guo2023animatediff, ), Make Pixel Dance (zeng2023make, ), PIA(zhang2023pia, ), and LivePhoto (chen2023livephoto, ), represent the other prevalent approaches. These models typically require inputting the entire video sequence during training and employ motion modules (guo2023animatediff, ) for temporal and positional encoding. However, limited by GPU memory constraints, most current works are restricted to generating 16 frames. Focusing on practical applications, this paper proposes LoopAnimate, a novel image-to-video generation method addresses the challenge of generating longer video sequences and enabling seamless transitions between the first and last frames, achieving a total output of 35 frames at once.
To address the challenge of object fidelity in image-to-video generation tasks, we propose the Multilevel Image representation and Textual semantics Decoupling Framework (MITDF). For long video generation, we introduce a Three-Stage Training Strategy to progressively increase frames while reducing fine-tuning modules, along with a TEMM that extends temporal positional encoding capabilities to 36 frames.
Within MITDF, we utilize a pretrained image-to-image diffusion model (sd-image-variations-diffusers, ) as the foundation. Input image information is incorporated into the generation process at both pixel-level and feature-level through conditioning and image embedding, respectively. Image representation and textual semantic embedding are injected into different stages of the diffusion model via cross-attention. Experiments demonstrate that excessive image embedding information will limit the motion dynamics of the generated videos, while the middle block contributes most significantly. Consequently, image representation embedding is injected only during the down-sampling process of the diffusion model, and textual semantic embedding is injected solely into the middle and up-sampling blocks. Additionally, we observe that this architecture effectively mitigates the influence of WebVid10M training data  (Bain21, ) watermarks during inference.
@@
  • Long Video Generation. We present a Three-Stage Training Strategy that progressively increases the number of generated frames in each stage while reducing the fine-tuning modules. Combined with the TEMM, which extends the temporal encoding length to 36 frames. our approach enables the generation of 35-frame videos in a single pass.
@@ 2.2. Long Video Generation
  • Stable Video Diffusion  (blattmann2023stable, ) presents a latent video diffusion model for high-resolution, state-of-the-art image-to-video generation. And SVD-XT can generate good quality image-conditioned videos with 25 frames in 2 to 8 steps with 576$\times$1024 resolutions.
@@ 3. Method
Figure 2. Illustration of LoopAnimate. Arrows for different training stages are with different colors.
@@
Multi-level image representation leverages information from both the pixel-level and feature-level of the input image. As illustrated in Figure. 2, during training, the first and last frames (marked as $F(0)$ and $F(-1)$) of the video are used as one of the conditions to the UNet in the form of Static and Dynamic Stacking Latent Code (SDSLC), enabling pixel-level information injection. Additionally, image embedding is extracted from the input image and cross-attention is performed during the down-sampling process, encoding the image representation information into the latent space and achieving feature-level information injection.
Appearance-semantic decoupling involves injecting the image representation embedding and text semantic embedding at different layers of the denoising network. Inspired by  (basu2023localizing, ), which suggests that different layers of the denoising network in text-to-image tasks have varying influences on style, objects, color, and motion, we conducted ablation experiments in Sec. 4.2.1. Similar to text-to-image tasks, we found that the middle block has the most significant impact on motion. We further discovered that excessive insertion of image embedding will restrict the motion dynamics of the generated video. Based on these findings, we finalized the network architecture as shown in Figure. 2. Additionally, we increased the temporal position encoding length in the temporal attention module within the motion module, naming it the Temporal Enhanced Motion Module. We were pleasantly surprised to find that this structure effectively avoids the influence of watermarks present in the WebVid10M training data  (Bain21, ) during inference.
In this section, we introduce the (1) fundamentals of diffusion models for image and video generation as a foundation. Subsequently, the proposed method is presented in detail with Figure 2, following the order of the (2) Asymmetric Loop Sampling Strategy for training data, the (3) specially designed condition initialization method and the (4) MITDF, and the (5) three-stage training strategy.
@@ 3.3.3. Fine-Grained Semantic Mask
During the inference stage, the object’s mask $M_{0}$ is obtained from the input image. We then create a tensor $Z_{M}$ by replicating one padding to all frames except for the first and last frames ${Z}_{M}={[{M}_{0},OnePAD,{M}_{0}]}$. Finally, $Z_{t}$, $Z_{SD}$, and $Z_{M}$ are concatenated along the channel dimension to form the final input condition ${Z}_{C}$ for the UNet, denoted as Equation (11), where ${{Z}_{C}}\in\mathcal{R}^{n*9*(2f-1)*h*w}$,
@@
3.4. Multi-level Image representation and Textual semantics Decoupling Framework
Object fidelity is a crucial aspect of image-to-video generation. To address this challenge, we propose a framework built upon a pretrained image-to-image diffusion model which initially utilizes image embedding as the guidance, and we enable the capability of decoupling multi-level image appearances and textual semantics. This decoupling involves injecting image appearance embedding and text semantic embedding at different layers of the denoising network. Multi-level image representation refers to incorporating pixel-level representation information of the input image during condition initialization, see Equation (8). And extract image embedding from the input image to perform cross-attention to inject the feature-level information into the latent space. Our ablation experiments reveal that, similar to text-to-image tasks (rombach2022highresolution, ; pixart-alpha, ; Imagen, ; rapheal, ; qu2023layoutllm, ; sun2023sgdiff, ), text semantics in image-to-video tasks exert the most significant influence in the middle block. Through limited ablation experiments in Sec. 4.2.1, we ultimately determined the optimal structure representation to be Equation (12) and Equation (13).
@@ 3.4. Multi-level Image representation and Textual semantics Decoupling Framework
Similar to Equation (14), as shown in Equation (15), ${Q}_{Z}^{M,U}$ represents the query extracted from the latent code ${Z}$ of the middle and up-sample layers, while ${K}_{T}$ and $V_{T}$ denote the key and value obtained from the text embedding, respectively.
@@
${{W_{q}},{W_{k}},{W_{v}}}$ are the weight matrices of the trainable linear projection layers in down-sample blocks, and ${{W_{q}^{{}^{\prime}}},{W_{k}^{{}^{\prime}}},{W_{v}^{{}^{\prime}}}}$ are in middle and up-sample blocks. ${{c_{i}},{c_{t}}}$ are image embedding and text embedding separately.
@@ 3.5. Three-Stage Training Strategy
Existing UNet-based video generation models require the input of an entire video for temporal and positional information encoding. However, limited by hardware memory, most existing video generation works can only generate $16$ frames at a time. To address this issue, we propose a three-stage training strategy which progressively increases the number of generated frames while reducing the fine-tuning modules in each stage. Additionally, we introduce the TEMM, which extends the temporal and positional information encoding length to $36$ frames.
@@
Stage2. When the input image contains a salient object, maintaining object consistency and action continuity in the generated results becomes crucial, increasing the difficulty of this specific type of video generation task. To tackle this challenge, we leverage SOD  (liu2024lightweight, ) to determine whether the first frame of the video contains the same object as described in the caption. To identify whether the object exists within a video, we employed optical flow  (teed2020raft, ). This process allows us to curate the Salient dataset, which contains 94,686 high-quality video segments. In the second stage, we train conv in and TEMM with $f=11$ ($21$ frames). Since the change of $f$ alters the encoding length of the positional embedding, TEMM needs to be fine-tuned. Additionally, fine-tuning conv in increases negligible computational cost but can accelerate convergence.
@@ 4. Experiments
Stage 2. The model is fine-tuned on a processed dataset of 94,686 high-quality videos containing salient objects (Salient Dataset). The conv in layer and TEMM are fine-tuned with $f=11$, enabling the generation of 21 frame videos. The learning rate is fixed at $56.e-$ 5, and the fine-tuning process runs for 100k iterations.
Stage 3. The model is further fine-tuned on the Salient Dataset, specifically focusing on conv in and the Q and V components of the temporal attention module within the TEMM. We increase $f$ to 18, enabling the generation of 35-frame videos. The learning rate is set to $52.e-$ 5, and the fine-tuning process runs for another 100k iterations. Table 1 shows the number of parameters fine-tuned in each stage. In Stage 3, only the conv in layer, TEMM.Q, and TEMM.V are fine-tuned, totaling $303.2M$ parameters. The number of frames selected for the three-stage training process is flexible. However, an excessively large number of frames in the first stage can prolong the training time. The temporal positional encoding length in TEMM can be extended, as long as it doesn’t exceed the GPU memory limit during the third stage fine-tuning. In our tests, using an A100 GPU with 80GB memory to run a single batch without any engineering optimization, the memory limit was reached when the third stage fine-tuning length was around 35 frames.
@@ 4.1. Comparison Results
We benchmarked our LoopAnimate against state-of-the-art I2V methods using objective metrics and subjective visual comparisons.
4.1.1. Quantitative Results
We selected three open-source I2V methods, I2VGen-XL  (zhang2023i2vgenxl, ), PIA (zhang2023pia, ), and SVD-XT  (blattmann2023stable, ), as our baselines for comparison. We conducted experiments on a test set consisting of 99 images, and the results are presented in Table 2. In Stage 3, only the conv in, TEMM.Q, and TEMM.V modules were fine-tuned, achieved the best results across all metrics except Loop-C, while still being able to generate 35 video frames at once. The superior performance in CLIP-I and ${MSE}_{F_{0}}$ metrics indicates that our  LoopAnimate generates videos with high object fidelity, supporting the second point in our contributions regarding the effectiveness of the multi-level image appearance and textual semantic decoupling framework. Furthermore, the three-stage fine-tuning strategy successfully extends the generation length to 35 frames while maintaining video quality, validating the third point in our contributions.
@@ 4.1.1. Quantitative Results
I2VGen-XL (zhang2023i2vgenxl, ) 0.882 54.0 0.972 0.740 0.812 32
@@ 4.1.2. Visualization Results
Objective metrics fail to accurately assess the subjective visual quality of generated videos. Therefore, we opted for a subjective visual comparison experiment, employing several currently available methods: the open-source I2VGen-XL (zhang2023i2vgenxl, ), PIA (zhang2023pia, ), SVD-XT (blattmann2023stable, ) and Dynamicrafter (xing2023dynamicrafter, ), alongside the commercially available Runway Gen2  (gen2, ) and Pika1.0  (pikalab, ). The comparison results, as illustrated in Figure. 3, demonstrate that our proposed LoopAnimate method exhibits superior performance in both object fidelity and motion quality, particularly in the challenging task of generating real human portraits.
Furthermore, we conducted a user study involving 19 participants. Considering the extensive task volume, we selected 20 videos from a pool of 99 (as mentioned in Sec.  4.1.1) and evaluated the subjective quality based on two aspects: subject consistency ${Q}_{S}$ and motion quality ${Q}_{M}$. The scores ranged from 1 to 3, with 3 being the highest. The statistical results are presented in Table 3. Our LoopAnimate method achieved the highest scores for both ${Q}_{S}$ and ${Q}_{M}$. It should be noted that SVD-XT  (blattmann2023stable, ) lacks text input. Consequently, its ${Q}_{M}$ score is significantly lower due to the evaluation’s consideration of consistency with the provided textual prompts during the subjective assessment.